# Visual Position Prompt for MLLM based Visual Grounding

### v1

@@ I Introduction
While research [20, 21] indicates that MLLMs possess a reasonable ability for spatial understanding, there is a growing consensus that these models still require further enhancement, particularly for tasks involving precise spatial reasoning like visual grounding. As illustrated in Fig. 1 (a), the grounding results from LLaVA-v1.5, for instance, reveal notable inaccuracies (the red box indicates the predicted bounding box, while the green box represents the ground truth). Although the predicted bounding box partially covers the target object, i.e., the brown toy, it suffers from both size and shape inaccuracies, failing to align well with the object’s true boundaries. These limitations highlight the need for more effective strategies to improve spatial alignment and object localization in MLLMs. To address these challenges, some studies are investigating the integration of advanced region-level enhancement modules and larger, more comprehensive visual grounding datasets into MLLMs [13, 22, 23, 6]. Other approaches are exploring the incorporation of task-specific expertise, such as converting special tokens directly into bounding boxes with additional decoder, to improve localization accuracy [24, 25].
However, despite the use of the aforementioned methods to enhance MLLM performance in grounding tasks, studies suggest that MLLMs still face significant challenges in precisely aligning coordinates with spatial information in images [26, 27]. One key issue lies in the models’ ability to effectively interpret and utilize spatial cues, which remain underutilized in many cases. As shown in Fig. 1 (b), when we introduce a positional reference in the form of a coordinate axis, providing an explicit spatial guide, the model’s understanding of spatial relationships improves significantly. This reference allows the model to better interpret the spatial information of objects in the image, leading to more accurate localization. Specifically, the predicted bounding box becomes more aligned with the brown toy, reflecting the improved spatial reasoning and localization accuracy when positional references are incorporated.
@@ II-A Visual Grounding
MLLM-based methods. With the tremendous success of LLMs in the field of natural language processing [40, 41], researchers are considering expanding their application to the multimodal domain [2, 3]. LLaVA [2] and MiniGPT-4 [42] are typical examples of such efforts. However, they still have room for improvement in the field of visual grounding. To enhance MLLMs’ grounding capability, researchers are proposing the incorporation of more visual grounding data and advanced position-aware modules. For example, Shikra [20], Kosmos [43], and MiniGPT-v2 [13] discrete coordinates into specialized tokens to better align with MLLMs’ input formats. Ferret [7] introduces a Spatial-Aware Visual Sampler, while PINK [44] employs a self-consistent bootstrapping approach to enhance region sensitivity. These methods still adhere to the standard MLLM framework, where the decoder of the MLLM directly outputs the coordinates. Since LLMs are inherently not well-suited for dense prediction tasks, another type of MLLM-based grounding method exists, which focuses on introducing a task-specific decoder. For example, LISA [24] and GLaMM [45] leverages the SAM decoder for pixel-level referring image segmentation while LLaVA-grounding [25] introduces an additional grounding module to predict bounding boxes. Although they have achieved great progress in visual grounding, they still face challenges in precisely aligning coordinates with spatial information in images. Besides, almost all these models are trained on large-scale datasets. For example, MiniGPT-v2 [13] trains on approximately 21M grounding data samples. In contrast, our approach explicitly aids MLLMs in learning these correspondences by providing clear positional references, achieving state-of-the-art performance while requiring training on only a much smaller dataset.
@@ IV-C Main Results
Finally, RefCOCOg, which contains long and more complex queries, is an even more challenging dataset. Despite differences in data scale and visual encoder capabilities, our method achieves competitive performance on RefCOCOg compared to current state-of-the-art methods such as Qwen-VL-7B and Groma. This phenomenon further demonstrates the effectiveness of our proposed method. By incorporating VPPs, our MLLM can more effectively align spatial information within images to coordinates, even when trained on significantly smaller datasets (0.6M vs. over 20M).

### v2

@@ I Introduction
While research [20, 21] indicates that MLLMs possess a reasonable ability for spatial understanding, there is a growing consensus that these models still require further enhancement, particularly for tasks involving precise spatial reasoning like visual grounding. As illustrated in Fig. 1 (a), the grounding results from LLaVA-v1.5, for instance, reveal notable inaccuracies (the red box indicates the predicted bounding box, while the green box represents the ground truth). Although the predicted bounding box partially covers the target object, i.e., the brown toy, it suffers from both size and shape inaccuracies, failing to align well with the object’s true boundaries. These limitations highlight the need for more effective strategies to improve spatial alignment and object localization in MLLMs. To address these challenges, some studies are investigating the integration of advanced region-level enhancement modules and larger, more comprehensive visual grounding datasets into MLLMs [13, 22, 23, 6, 24]. Other approaches are exploring the incorporation of task-specific expertise, such as converting special tokens directly into bounding boxes with additional decoder, to improve localization accuracy [25, 26].
However, despite the use of the aforementioned methods to enhance the performance of MLLM in grounding tasks, studies suggest that MLLMs still face significant challenges in precisely aligning coordinates with spatial information in images [27, 28]. One key issue lies in the models’ ability to effectively interpret and utilize spatial cues, which remain underutilized in many cases. As shown in Fig. 1 (b), when we introduce a positional reference in the form of a coordinate axis, providing an explicit spatial guide, the model’s understanding of spatial relationships improves significantly. This reference allows the model to better interpret the spatial information of objects in the image, leading to more accurate localization. Specifically, the predicted bounding box becomes more aligned with the brown toy, reflecting the improved spatial reasoning and localization accuracy when positional references are incorporated.
@@ II-A Visual Grounding
MLLM-based methods. With the tremendous success of LLMs in the field of natural language processing [41, 42], researchers are considering expanding their application to the multimodal domain [2, 3]. LLaVA [2] and MiniGPT-4 [43] are typical examples of such efforts. However, they still have room for improvement in the field of visual grounding. To enhance MLLMs’ grounding capability, researchers are proposing the incorporation of more visual grounding data and advanced position-aware modules. For example, Shikra [20], Kosmos [44], and MiniGPT-v2 [13] discrete coordinates into specialized tokens to better align with MLLMs’ input formats. Ferret [7] introduces a Spatial-Aware Visual Sampler, while PINK [45] employs a self-consistent bootstrapping approach to enhance region sensitivity. These methods still adhere to the standard MLLM framework, where the decoder of the MLLM directly outputs the coordinates. Since LLMs are inherently not well-suited for dense prediction tasks, another type of MLLM-based grounding method exists, which focuses on introducing a task-specific decoder. For example, LISA [25] and GLaMM [46] leverage the SAM decoder for pixel-level reference image segmentation, while LLaVA grounding [26] introduces an additional grounding module to predict bounding boxes. Although they have achieved great progress in visual grounding, they still face challenges in precisely aligning coordinates with spatial information in images. In addition, almost all of these models are trained on large-scale datasets [13, 24, 6, 20]. For example, MiniGPT-v2 [13] trains on approximately 21M grounding data samples. In contrast, our approach explicitly aids MLLMs in learning these correspondences by providing clear positional references, achieving state-of-the-art performance while requiring training on only a much smaller dataset.
@@ IV-C Main Results
Finally, RefCOCOg, which contains long and more complex queries, is an even more challenging dataset. Models with relatively small amounts of training data, such as Shikra and GroundingGPT, perform less well. Fortunately, despite differences in data scale and visual encoder capabilities, our method achieves competitive performance on RefCOCOg compared to current state-of-the-art methods such as Qwen-VL-7B and Groma. This phenomenon further demonstrates the effectiveness of our proposed method. By incorporating VPPs, our MLLM can more effectively align spatial information within images to coordinates, even when trained on significantly smaller datasets (0.6M vs. over 20M).